home excer.

The partnership of King, Queen and Page has experienced operating losses for three consecutive years. The partners, who shares profits and losses in the ratios of King of 15%; Queen of 60%; and Page of 25%, are liquidating the business. The following is the condensed balance at December 31, 2018.

Cash                            $ 7,000                       Accounts Payable                   $63,000

Noncash Assets          163,000                       King, Capital                          24,000

                                                                        Queen, Capital                                    66,000

                                                                        Page, Capital                          17,000

Total Assets                170,000                       Total Liabilities and Capital 170,000

Requirements:

Prepare a liquidation (cash payments) schedule assuming the noncash assets are sold for $103,000.
Answer #1

Please find below table useful to compute desired results: -

A B C D E E F G H queen 66000 =-D4-B5)*60% =J4+J5 page 17000 =-(D4-B5) 25% =L4+L5 2 3 Items Cash 4 Balance Before liquidation

End results would be as follows: -

B C D E F G H I J K L 1 Liabilities 63,000 + king | + 24.000 (9,000) 15,000 queen | + 66,000 (36.000) 30,000 page 17,000 (15,